When solutions containing silver ions and chloride ions are mixed, silver chloride precipitates: Ag+(aq)+Cl−(aq)→AgCl(s)ΔH=−65.5kJ

Calculate ΔH for formation of 0.490 mol of AgCl by this reaction.

Calculate ΔH for the formation of 7.50 g of AgCl.

Calculate ΔH when 9.23×10−4 mol of AgCl dissolves in water.
Answer #1

The reaction is

Ag+(aq) + Cl-(aq) → AgCl(s) ΔH = −65.5kJ

for 1 mole of AgCl   ΔH = −65.5kJ then for

ΔH for formation of 0.490 mol of AgCl

(0.490)\times(-65.5) = -32.095 KJ

ΔH for the formation of 7.50 g of AgCl.

molar mass of AgCl = 143.32g/mol

7.50 gm = 7.50/143.32 = 0.05233 mole

(0.05233)\times(-65.5) = -3.4276 KJ

ΔH when 9.23×10-4 mol of AgCl dissolves in water.

rection is reversed thus,

AgCl(s) → Ag+(aq) + Cl-(aq) ΔH = +65.5kJ

(9.23×10-4)\times(65.5) = 0.06045 KJ